### v5

# MiLe Loss: a New Loss for Mitigating the Bias of Learning Difficulties in Generative Language Models

Abstract
Generative language models are usually pretrained on large text corpus via predicting the next token (i.e., sub-word/word/phrase) given the previous ones. Recent works have demonstrated the impressive performance of large generative language models on downstream tasks. However, existing generative language models generally neglect an inherent challenge in text corpus during training, i.e., the imbalance between frequent tokens and infrequent ones. It can lead a language model to be dominated by common and easy-to-learn tokens, thereby overlooking the infrequent and difficult-to-learn ones. To alleviate that, we propose a MiLe Loss function for mitigating the bias of learning difficulties with tokens. During training, it can dynamically assess the learning difficulty of a to-be-learned token, according to the information entropy of the corresponding predicted probability distribution over the vocabulary. Then it scales the training loss adaptively, trying to lead the model to focus more on the difficult-to-learn tokens. On the Pile dataset, we train generative language models at different scales of 468M, 1.2B, and 6.7B parameters. Experiments reveal that models incorporating the proposed MiLe Loss can gain consistent performance improvement on downstream benchmarks.
1 Introduction
Generative language models like GPT-3 Brown et al. (2020) are generally pretrained on extensive textual data, in the manner of predicting the next token given the previous ones for each training text. Recently, large generative language models have been exhibiting impressive performance on various downstream natural language tasks, like dialogue system, classification, sequence labeling, etc. Touvron et al. (2023); Brown et al. (2020); Chowdhery et al. (2022), and attracting much attention from both academia and industry.
Frequency Bucket high medium low
PPL 4.323 13.541 15.517
Table 1: The average perplexity (PPL) for tokens in different frequency buckets.
However, previous works have overlooked an inherent issue in natural language corpus that might affect the pretraining of a language model, i.e., frequent tokens far outnumber infrequent ones. Actually, Zipf’s law Piantadosi (2014) highlights the inherent imbalance of tokens in natural language datasets, i.e., a few frequent tokens would dominate a dataset while many infrequent ones only form a minor portion. For instance, $50\%$ of the Brown Corpus  Francis and Kucera (1979), which comprises over a million tokens, is covered by only the top 135 most frequent tokens.
The imbalance of tokens is essentially a class imbalance problem. We argue that infrequent tokens are difficult to learn due to their fewer occurrences, in contrast to the frequent ones that can be learned adequately Lin et al. (2017). To confirm that, we utilize the remarkable language model LLaMA Touvron et al. (2023) with 6.7B parameters on the Pile Gao et al. (2021a) validation set and perform a detailed perplexity (PPL) analysis at the token level. It’s worth noting that a higher perplexity is indicative of a token’s higher learning difficulty. In our analysis, all tokens are grouped into three frequency buckets: high, medium, and low, based on their counts in the whole Pile dataset [^1] As the Pile dataset is large enough, the relative frequencies of all tokens are supposed to be almost the same as those in the training set of LLaMA, which is not publicly available.. Here, we calculate the frequency of each token and sort them in descending order of frequency. Then, we categorize the top tokens that cover $80\%$ of the dataset as tokens of high frequency, those that cover the extra $15\%$ (i.e., $80\%-95\%$) of the dataset as tokens of medium frequency, and the remaining $5\%$ as tokens of low frequency. As shown in Table 1, for the tokens of high frequency, LLaMA derives a much lower average perplexity ($4.394$) than those of medium ($13.891$) or low ($15.814$) frequency. That confirms our assumption: token imbalance can lead to the bias of learning difficulties. More explicitly, those frequent and easy-to-learn tokens (i.e., classes) might overwhelm the model and make it neglect the infrequent and difficult-to-learn ones during training Lin et al. (2017). Therefore, we emphasize that the latter kinds of tokens should be given more attention during language model pretraining.
Figure 1: An example where predicting the next token is more like a multi-label classification problem.
It is a straightforward idea to use the notable Focal Loss Lin et al. (2017) from the field of object detection as an alternative to the prevalent Cross-Entropy Loss for the next token prediction. This modification aims to intensify the language model’s focus on the infrequent and difficult-to-learn tokens. Focal Loss is a dynamically scaled version of the Cross-Entropy Loss, where the scaling factor decreases as the predicted probability w.r.t the ground-truth token increases. Specifically, Focal Loss decreases the weights of the easy-to-learn tokens, as their predicted probabilities are higher, and meanwhile increases the weights of the difficult-to-learn ones, as their predicted probabilities are lower. In that way, it compels the language model to pay more attention to difficult-to-learn tokens.
Nevertheless, Focal Loss Lin et al. (2017) only takes into account the probability w.r.t the ground-truth token when assessing its learning difficulty, and is intuitively designed for the multi-class classification problem where an object is only associated with a single class label. Indeed, in language model pretraining, when predicting the next token given the previous ones, there might exist multiple valid tokens besides the ground-truth one. This makes predicting the next token more like a multi-label classification problem, where an object can be associated with multiple class labels. For example, as shown in Figure 1, given the previous tokens “I like playing ”, there are multiple valid next tokens, like “basketball”, “football”, “golf”, etc. Suppose the target training token sequence is “I like playing basketball”. As the valid tokens would divide up almost the total probability (i.e., 1.0), the ground-truth token “basketball” would be given a smaller probability (e.g., 0.18). Then Focal Loss would treat “basketball” for the position as a difficult-to-learn token. However, as all the other valid tokens are also correct for the position in the view of language modeling, only allowing “basketball” to be predicted is unsuitable. Thus, the learning difficulty assessed by Focal Loss for “basketball” is imperfect in such a multi-label classification case.
In this paper, we propose a new loss function termed MiLe Loss to better enable a language model to pay more attention to the difficult-to-learn tokens in such multi-label classification cases. We observe that when a next target token is easy-to-learn, the minor valid tokens would divide up almost the total probability while others are associated with very low probabilities, resulting in a low information entropy of the predicted probability distribution over the vocabulary. On the contrary, if a next token is difficult-to-learn, the predicted probability distribution would be more uniform, resulting in a higher information entropy. Therefore, instead of relying on the single probability of the ground-truth token as Focal Loss, the proposed MiLe Loss uses the information entropy of the predicted probability distribution for assessing learning difficulties, which can better handle cases with multiple valid tokens. Then, tokens exhibiting high-entropy, possibly being difficult-to-learn, will be assigned increased weights during language model pretraining.
To validate the effectiveness of the proposed MiLe Loss, we train three different-sized models on the Pile dataset Gao et al. (2021a). Experimental results indicate that MiLe Loss steadily outperforms Focal Loss and Cross-Entropy Loss on downstream benchmarks.
Our contributions can be summarized as follows.
  • We highlight the bias of learning difficulties in generative language models, which is mainly caused by the inherent token imbalance in textual training data.
  • We propose a new loss function termed MiLe Loss to enhance Focal Loss for mitigating the bias of learning difficulties.
  • We validate the effectiveness of the proposed MiLe Loss with extensive experiments. Experimental results show that it consistently outperforms Focal Loss and Cross-Entropy Loss.
2 Related Works
2.1 Language Models
Language Models are statistical models that aim to maximize the likelihood of the training sequences of tokens Touvron et al. (2023). Early language models are based on the statistics of $n$-grams Bahl et al. (1983); Katz (1987); Kneser and Ney (1995). Then the focus has shifted toward neural-network-based models. Recurrent Neural Networks Mikolov et al. (2010) and their variants, e.g., LSTMs Graves (2013), have been successful in this regard. Those models are capable of learning complex patterns in textual data and have achieved remarkable results in various language modeling tasks.
Recently, Transformers are commonly used as the backbone network for language models. Representative works include BERT Devlin et al. (2019), RoBERTa Liu et al. (2019), GPT-2 Radford et al. (2019), UniLM Dong et al. (2019a), and T5 Raffel et al. (2020), etc. Since the advent of GPT-3 Brown et al. (2020) with 175 billion parameters, which achieves outstanding performance in various downstream tasks, the research landscape has increasingly pivoted towards large generative language models. Notable works like Gopher Rae et al. (2021), Pythia Biderman et al. (2023), PaLM Chowdhery et al. (2022), GaLM Du et al. (2022), OPT Zhang et al. (2022) and LLaMA Touvron et al. (2023), have also been proposed.
However, previous works do not consider the bias of learning difficulties among tokens, which is mainly caused by the inherent token imbalance in the textual training data. They probably overlook some difficult-to-learn but informative tokens during model training. To tackle that, in this paper we introduce MiLe Loss, aiming to lead generative language models to pay more attention to those tokens.
2.2 Class Imbalance
Class Imbalance refers to a highly skewed distribution of classes in the training data, which means that the number of instances in some classes is significantly higher than those in the other classes Yang and Xu (2020). A commonly used solution is to perform data re-sampling, where the minority classes are up-sampled Chawla et al. (2002); Ando and Huang (2017); Pouyanfar et al. (2018); Shen et al. (2016), and the majority classes are down-sampled Lee et al. (2016); Buda et al. (2018). Other works Cui et al. (2019); Dong et al. (2019b); Lin et al. (2017) have also proposed enhanced loss functions to mitigate issues caused by class imbalance, e.g., Focal Loss.
In language modeling, to mitigate the mentioned bias of learning difficulties caused by the inherent token imbalance, one may simply refer to the data re-sampling method. However, data re-sampling at the token level, i.e., up-sampling infrequent tokens and down-sampling frequent ones, will probably break the semantics of training texts. Meanwhile, re-sampling at the coarse-grained sentence/paragraph/document/domain level will equally increase/decrease the number of both kinds of tokens, and thus cannot well tackle the token imbalance.
Therefore, we consider enhancing the loss function to alleviate the bias of learning difficulties among tokens for generative language models, enabling them to pay more attention to those difficult-to-learn but informative tokens. Firstly, we attempted to use the notable Focal Loss. However, since predicting the next token in generative language models is more like a multi-label classification problem as analyzed before, Focal Loss struggles to give suitable scaling factors for cases with multiple valid next tokens. To tackle that, we introduce the MiLe Loss.
3 Method
3.1 Preliminaries
Language Model Pretraining
As mentioned before, a generative language model is generally trained via predicting the next token (i.e., sub-word/word/phrase), one by one, based on the previous ones for each training text, aiming to maximize the likelihood. Formally, given a training text $T$ consisting of $n$ tokens, i.e., $T=[t_{1},\ldots,t_{i-1},t_{i},\ldots,t_{n}]$, when predicting a target token $t_{i}$, the generative language model takes the previous ones $\mathbf{t}=[t_{1},t_{2},...,t_{i-1}]$ as input, and then generates a probability distribution $\mathbf{p}$ over the vocabulary as output. In nearly all implementations, the Cross-Entropy loss is employed as the loss function, to maximize the predicted probability $\mathbf{p}_{t_{i}}$ w.r.t the ground-truth token $t_{i}$. Considering that the recent state-of-the-art deep language models (LM) predominantly leverage the Transformer architecture Vaswani et al. (2017), the training loss $\mathcal{L}_{CE}$ of the generative language model can be formulated as follows.
$\mathcal{L}_{CE}$ $=-\log(\mathbf{p}_{t_{i}})$ (1)
$s.t.,\quad\mathbf{p}$ $=\text{softmax}(W\mathbf{H}_{i-1}^{last})$ (2)
$\mathbf{H}^{last}$ $=\text{Transformer}(\text{Embedding}(\mathbf{t}))$ (3)
Here, $\mathbf{H}^{last}$ denotes the hidden states of the last layer of the Transformer architecture, which consists of the hidden states w.r.t the previous tokens $\mathbf{t}=[t_{1},t_{2},...,t_{i-1}]$, i.e., $\mathbf{H}^{last}=[\mathbf{H}_{1}^{last},\mathbf{H}_{2}^{last},\ldots,\mathbf{H}_{i-1}^{last}]$. With $\mathbf{H}^{last}_{i-1}$, a linear projection layer $W$ is introduced to derive the predicted probability distribution $\mathbf{p}$ over the vocabulary, with a softmax operation.
Focal Loss for Classification
Focal Loss is originally proposed for object detection to address the issue of extreme foreground-background class imbalance encountered during the training of one-stage object detectors Lin et al. (2017). Focal Loss can lead a classification model to concentrate more on a sparse set of difficult-to-learn classes and prevent the abundance of easy-to-learn classes from overwhelming the model during training. Actually, Focal Loss is an extension of Cross-Entropy Loss, with an extra dynamic scaling factor, as formulated below.
$\mathcal{L}^{0}_{FL}$ $=-(1-p)^{\gamma}\log(p)$ (4)
Here, $p$ is the predicted probability w.r.t the ground-truth class, and $\gamma$ is a hyperparameter with $\gamma\geq 0$. It can be seen that when $\gamma=0$, Focal Loss would degenerate to Cross-Entropy Loss. As $p$ decreases, i.e., getting more-difficult-to-learn, the dynamic scaling factor $(1-p)^{\gamma}$ increases, thus giving more attention (i.e., higher weights) to the difficult-to-learn classes.
3.2 Focal Loss for Language Models
Generative language models are commonly trained on the massive textual corpus, which exhibits inherent token imbalance as revealed by Zipf’s law Piantadosi (2014). Such an imbalance of tokens can lead to two primary challenges: 1) Training efficiency becomes sub-optimal. A large number of easy-to-learn tokens (i.e., classes) provide marginal gains in learning signals. Lin et al. (2017). 2) The training process can be overwhelmed by a large proportion of the frequent and easy-to-learn tokens, and thus pay insufficient attention to the other infrequent, difficult-to-learn but informative tokens, which might lead to performance degradation.
As revealed in Equation (1), training a generative language model is essentially a classification problem. Therefore to mitigate the bias of learning difficulties caused by the inherent token imbalance, Focal Loss can be applied. Specifically, we can use the Focal Loss as a substitute for the Cross-Entropy Loss in Equation (1) to train a generative language model as follows.
$\mathcal{L}_{FL}$ $=-(1-\mathbf{p}_{t_{i}})^{\gamma}\log(\mathbf{p}_{t_{i}})$ (5)
Here, the dynamic scaling factor $(1-\mathbf{p}_{t_{i}})^{\gamma}$ is derived based on the predicted probability $\mathbf{p}_{t_{i}}$ of the to-be-learned token $t_{i}$. Similarly, as the probability $\mathbf{p}_{t_{i}}$ decreases (i.e., being more difficult to learn), the scaling factor $(1-\mathbf{p}_{t_{i}})^{\gamma}$ increases correspondingly. Therefore, more-difficult-to-learn tokens will receive higher loss weights.
3.3 Proposed MiLe Loss
However, as illustrated in Figure 1 and analyzed before, in language model pretraining, predicting the next token is more like a multi-label classification problem. When there are multiple valid next tokens for a given sequence of previous tokens, the learning difficulty assessed by Focal Loss is imperfect.
To tackle that, we propose MiLe Loss, which leverages the information entropy of the predicted probability distribution $\mathbf{p}$ over the vocabulary, instead of the single probability $\mathbf{p}_{t_{i}}$ as Focal Loss, to derive a dynamic scaling factor. MiLe Loss is naturally designed for cases with multiple valid tokens. It is inspired by the following observations: 1) when a next token is easy-to-learn, the minor valid tokens would divide up almost the total probability (i.e., 1.0) while others are associated with very low probabilities (i.e., $\mathbf{p}$ is more focused), resulting in a low information entropy; 2) when a next token is difficult-to-learn, the predicted probability distribution would be more uniform, resulting in a higher information entropy.
model size dimension $n$ heads $n$ layers learning rate batch size seq length
468M 1024 16 24 $3.0e^{-4}$ 1024 1024
1.2B 2048 8 16 $3.0e^{-4}$ 1024 1024
6.7B 4096 32 32 $3.0e^{-4}$ 2048 2048
Table 2: Model sizes, architectures, and optimization hyper-parameters.
Specifically, MiLe Loss can be formulated as follows in language model pretraining.
$\mathcal{L}_{IL}$ $=-(1-\sum_{j}\mathbf{p}_{j}\log(\mathbf{p}_{j}))^{\gamma}\log(\mathbf{p}_{t_{i}})$ (6)
Here, $-\sum_{j}\mathbf{p}_{j}\log(\mathbf{p}_{j})\geq 0$ is the information entropy of the predicted probability distribution $\mathbf{p}$ over the vocabulary. Note that when $\mathbf{p}$ is a uniform distribution, i.e., $p_{j}=\frac{1}{N}$ with $N$ being the vocabulary size for all $j$, the information entropy reaches its upper bound $\log(N)$. Therefore, the dynamic scaling factor $(1-\sum_{j}\mathbf{p}_{j}\log(\mathbf{p}_{j}))$ is bounded in $[1,1+\log(N)]$. When a next token is difficult to learn, the corresponding higher information entropy results in a higher scaling factor, and thus MiLe Loss increases the loss weights for such tokens. Conversely, MiLe Loss decreases the loss weights for easy-to-learn tokens, according to their lower information entropies.
4 Experiments
We train three generative language models of different capacities, i.e., 468M, 1.2B, and 6.7B parameters, on the open-source Pile dataset  Gao et al. (2021a) as  Biderman et al. (2023); Xie et al. (2023); Carlini et al. (2023), and make comparisons among different loss functions.
4.1 The Pile dataset
The Pile dataset is a public large-scale corpus for language model pretraining, which has over 825GB English texts across 22 domains. For experiments, we tokenize it using the remarkable LLaMA tokenizer  Touvron et al. (2023) with a 32k-sized vocabulary. As the number of tokens changes with a new tokenizer, we follow Xie et al. (2023) to re-calculate the sampling weight for each domain. Specifically, we chunk the dataset into sequences of 1,024 tokens, and then for each domain, we multiply its corresponding number of sequences with its domain-specific epochs reported in Gao et al. (2021a). Finally, we normalize all the multiplication results to obtain the sampling weights listed in Table 3.
Weights Weights
ArXiv 0.1997 OpenSubtitles 0.0239
BookCorpus2 0.0100 OpenWebText2 0.1735
Books3 0.1640 PhilPapers 0.0073
DM Mathematics 0.0502 Pile-CC 0.1551
Enron Emails 0.0030 PubMed Abstracts 0.0536
EuroParl 0.0156 PubMed Central 0.2823
FreeLaw 0.0895 StackExchange 0.1027
Github 0.0962 USPTO Backgrounds 0.0586
Gutenberg(PG-19) 0.0481 Ubuntu IRC 0.0229
HackerNews 0.0117 Wikipedia(en) 0.1121
NIH ExPorter 0.0047 YoutubeSubtitles 0.0151
Table 3: Sampling weights on the Pile dataset.
4.2 Experimental setup
BoolQ
HellaSwag
LAMBADA
OpenBookQA
PIQA
SIQA
StoryCloze
Winogrande
Avg
468M
0-shot Cross-Entropy Loss
57.52
40.73
39.10
30.60
67.08
40.79
63.55
53.75
49.14
Focal Loss
58.35
41.17
40.09
32.80
67.25
41.91
63.07
51.70
49.54
MiLe Loss 59.57 41.27 41.34
30.00
67.25
41.61
63.60 54.78 49.93
1-shot Cross-Entropy Loss
54.22
40.86
37.16
30.40
67.85
41.66
62.69
53.04
48.48
Focal Loss
53.64
41.04
37.88
32.20
67.14
44.27
62.16
52.64
48.87
MiLe Loss 55.23
40.90
38.75
32.00
67.68
43.35
63.23 55.88 49.63
5-shot Cross-Entropy Loss
50.89
41.06
36.27
28.80
67.68
43.39
62.37
50.99
47.68
Focal Loss
48.10
41.80
38.50
31.40
67.19
46.01 63.01
52.09
48.51
MiLe Loss 52.29
41.53
39.05
28.80
67.41
45.39
62.85
54.06 48.92
1.2B
0-shot Cross-Entropy Loss
55.96
47.48
45.76
32.20
69.64
42.43
65.47
54.54
51.69
Focal Loss 62.02
47.61
46.87
33.00
69.59
42.02
65.63
55.01
52.72
MiLe Loss
56.94
47.64 47.37 33.80 70.13
41.91
66.06 55.96
52.48
1-shot Cross-Entropy Loss
54.71
47.37
42.13
34.40
69.42
44.78
65.26
56.27
51.79
Focal Loss 62.35 47.41
43.88
32.60
69.15
45.04
65.42
54.85
52.59
MiLe Loss
54.95
47.39
45.08 34.00 70.13 45.04 65.58
54.85
52.13
5-shot Cross-Entropy Loss
55.72
47.74
41.55
33.00
69.86
45.04
66.11
55.64
51.83
Focal Loss 62.17 48.00
42.87
32.00
69.75
45.60
66.01
56.20
52.82
MiLe Loss
55.38
47.78
45.00 34.00 70.13 46.26 66.22 56.83
52.70
6.7B
0-shot Cross-Entropy Loss 62.14
58.91
55.54
34.40
73.61
44.06
70.66
61.40
57.59
Focal Loss
59.72
59.59
55.64
36.60
73.94
43.04
70.12
61.88
57.57
MiLe Loss
60.89
59.63 57.73
35.20
73.99 44.06 71.25
61.01
57.97
1-shot Cross-Entropy Loss
59.24
58.68
53.48
37.00
73.99
47.90
70.60
60.69
57.70
Focal Loss
58.53
59.23
52.59
35.60
74.27
48.06
69.96
59.91
57.27
MiLe Loss 60.46 59.56 55.35 38.00
73.29
48.57 70.87 61.01 58.39
5-shot Cross-Entropy Loss
61.28
59.44
54.01
37.00
74.16
49.03
71.30
63.06
58.66
Focal Loss
57.98
60.10
55.91
36.80
74.05
50.0
70.44
62.90
58.52
MiLe Loss 62.20
60.06
58.16 37.80
73.61
50.67 71.67 63.30 59.68
Table 4: Zero-shot and few-shot performance (i.e., accuracy) of models at different scales on common sense reasoning benchmarks.
We train three generative language models with 468M, 1.2B, and 6.7B parameters, respectively. Specifically, the architectures of the 468M-parameter and the 1.2B-parameter models, including the dimensionality of hidden states, the number of layers, etc., are identical to those of the 410M-parameter and the 1.0B-parameter models outlined in Biderman et al. (2023). The minor differences in parameter sizes are attributed to the variations of vocabulary size in the embedding layer. As for the 6.7B-parameter model, its architecture is identical to LLaMA-7B Touvron et al. (2023). The corresponding hyperparameters for each model can be found in Table 2. Following LLaMA Touvron et al. (2023), we use the AdamW optimizer Loshchilov and Hutter (2019) with a learning rate of $3.0e^{-4}$, $2k$ warmup steps, and a cosine learning rate decay schedule. Following  Lin et al. (2017), the hyperparameter $\gamma$ is set as $1.0$ for both Focal Loss and the proposed MiLe Loss, unless explicitly stated otherwise. Due to the computational budget and following the pretraining settings of Xie et al. (2023), all models are pretrained with 100B tokens.
Following Touvron et al. (2023); Brown et al. (2020); Rae et al. (2021); Hoffmann et al. (2022), we primarily evaluate all models on tasks of common-sense reasoning, closed-book question answering, and massive multitask language understanding. For fair comparisons, we utilize the open-source pipeline lm-evaluation-harness [^2] https://github.com/EleutherAI/lm-evaluation-harness Gao et al. (2021b) for evaluation, as Biderman et al. (2023); Dettmers and Zettlemoyer (2023).
4.3 Experimental Results
Common Sense Reasoning
Following Touvron et al. (2023); Brown et al. (2020); Rae et al. (2021); Hoffmann et al. (2022), we employ 8 widely used benchmark datasets for the evaluation of common sense reasoning, including BoolQ Clark et al. (2019), HellaSwag Zellers et al. (2019), LAMBADA Paperno et al. (2016), OpenBookQA Mihaylov et al. (2018), PIQA Bisk et al. (2020), SIQA Sap et al. (2019), StoryCloze Mostafazadeh et al. (2016),Winogrande Sakaguchi et al. (2020). We report the model performance in terms of accuracy for zero-shot and few-shot settings in Table 4, like Touvron et al. (2023); Brown et al. (2020).
We can observe that the proposed MiLe Loss substantially outperforms both Cross-Entropy Loss and Focal Loss on different setups with different model capacities. Specifically, for models with 468M and 6.7B parameters on 0/1/5-shot settings, MiLe Loss consistently achieves superior performance to both compared baselines. As for the 1.2B-parameter model, although MiLe Loss yields slightly lower average performance than Focal Loss, it still delivers the highest performance on 6 out of the 8 datasets and steadily outperforms Cross-Entropy Loss on most datasets.
These results clearly demonstrate the effectiveness of the proposed MiLe Loss. We attribute it to that MiLe Loss compels language models to allocate more attention to those difficult-to-learn yet informative tokens during pretraining, which mitigates the bias of learning difficulties among tokens. Moreover, the consistent performance superiority of MiLe Loss over Focal Loss also validates that, relying on the information entropy of the predicted probability distribution over the vocabulary to assess the learning difficulties of tokens is more reasonable.
Closed Book Question Answering
0-shot 1-shot 5-shot
TriviaQA
Cross-Entropy Loss 17.09 21.98 26.33
Focal Loss 16.47 23.03 27.31
MiLe Loss 20.64 23.42 28.75
WebQuestions
Cross-Entropy Loss 5.22 9.79 14.17
Focal Loss 4.53 9.60 14.62
MiLe Loss 5.02 9.89 14.57
Table 5: Zero-shot and few-shot exact match performance of 6.7B-parameter models on closed-book question-answering benchmarks.
Following Brown et al. (2020); Touvron et al. (2023), for the task of closed book question answering, we evaluate the performance of the largest 6.7B-parameter models with different loss functions on two benchmark datasets, i.e., TriviaQA Joshi et al. (2017) and WebQuestions Berant et al. (2013). We report the exact match performance for the zero-shot and few-shot settings in Table 5.
It can be seen that language models trained with the proposed MiLe Loss achieve superior performance across most settings. Compared with Cross-Entropy Loss, MiLe Loss achieves substantial performance improvement in 5 out of 6 settings. Particularly, on TriviaQA, MiLe Loss achieves a maximum performance improvement of 3.55% (0-shot) over Cross-Entropy Loss. Compared with Focal Loss, MiLe Loss also exhibits consistent superiority. Notably, in the 0-shot setting on TriviaQA, MiLe Loss outperforms Focal Loss by 4.17%.
Massive Multitask Language Understanding
We further validate the effectiveness of the proposed MiLe Loss on the MMLU (Massive Multitask Language Understanding) benchmark Hendrycks et al. (2021). MMLU consists of multiple-choice questions covering 57 subjects, including STEM, social sciences, humanities, etc. It has been serving as a benchmark for evaluating the multitasking capability of pretrained language models. Following LLaMA Touvron et al. (2023), we evaluate the 6.7B-parameter models in the 5-shot setting. Among multiple choices, we choose the one with the highest probability normalized by the number of tokens.
As shown in Table 6, MiLe loss exhibits superior performance on average. Compared with Cross-Entropy loss, MiLe loss obtains performance improvement of 0.32%, 1.28%, and 0.91% for the field of STEM, Humanities, and Other, respectively. For the field of Social Sciences, the performance decline may be attributed to that MiLe Loss tends to consider Social Sciences samples as easier-to-learn ones. We intend to study it in depth in our future work. Compared with Focal Loss, MiLe Loss also yields superior performance on all fields except STEM. All the results above further demonstrate the proposed MiLe Loss’s effectiveness and reasonableness.
5 Analyses
We conduct further experiments to provide more insightful analyses on the proposed MiLe Loss.
5.1 Impact of $\gamma$
Cross-Entropy Focal MiLe
Loss Loss Loss
STEM 29.59 29.99 29.91
Social Sciences 29.64 27.57 28.07
Humanities 27.00 27.35 28.28
Other 29.94 29.34 30.85
Avg 29.38 28.90 29.68
Table 6: The 5-shot learning performance of 6.7B-parameter models on MMLU.
Figure 2: The performance of MiLe Loss and Cross-Entropy Loss in 5-shot learning with different $\gamma$ values.
We aim to discern the performance change of the proposed MiLe Loss on language models with different values of $\gamma$, i.e., the hyperparameter in Equation (6). It’s worth noting that when $\gamma$ is set to 0, MiLe Loss is functionally equivalent to Cross-Entropy Loss. As $\gamma$ increases, the language model becomes more focused on the difficult-to-learn tokens, i.e., those with higher information entropy. Here we conduct a grid search for $\gamma$ on language models of various scales (i.e., 468M, 1.2B, and 6.7B parameters), and use the average performance in 5-shot learning for the Common Sense Reasoning task that covers the most benchmarks as the evaluation metric.
As shown in Figure 2, when $\gamma$ increases from 00 to $5$ for the 468M-parameter model or increases from 00 to $2$ for the 1.2B-parameter/6.7B-parameter models, the performances of MiLe Loss consistently surpass those of Cross-Entropy Loss. The results clearly demonstrate that the performance of MiLe Loss is not very sensitive to the setting of the hyperparameter $\gamma$, which shows practical applicability. As expected, when $\gamma$ increases to a relatively large value, the performance of MiLe Loss declines, because too much attention is given to the difficult-to-learn tokens, and the easy-to-learn ones get overlooked as a result.
5.2 Perplexity on the Pile Validation Set
Here we further discuss how the proposed MiLe Loss affects the perplexity of pretrained language models on the Pile validation set.
Table 7 reports the perplexity of the largest 6.7B-parameter models trained with $\gamma$ increasing from 00 to $5$ for MiLe Loss. Among them, $\gamma=0$ is equivalent to Cross-Entropy Loss. Notably, when $\gamma=0.5$, the perplexity obtained by MiLe Loss is lower than that by Cross-Entropy Loss (i.e., $\gamma=0$). However, as we increase $\gamma$, the perplexity of MiLe Loss also increases and becomes higher than that of Cross-Entropy Loss. The increase of perplexity can be attributed to: 1) the measurement of perplexity is directly related to the exponentiation of Cross-Entropy Loss, and thus optimizing Cross-Entropy Loss during training is consistent with optimizing the perplexity; 2) the objective function of MiLe Loss somewhat diverges from that of perplexity due to the dynamic scaling factor, and thus optimizing it may lead to an increase of perplexity.
To thoroughly inspect how the perplexity increases, we conduct a fine-grained analysis of perplexity at the token level. Similar to the perplexity analysis before, we group all tokens into three learning-difficulty levels based on their corresponding frequencies, i.e., easy, medium, and difficult. Specifically, we categorize the top tokens that cover $80\%$ of the Pile dataset as easy, those that cover the extra $15\%$ (i.e., $80\%-95\%$) of the Pile dataset as medium, and the remaining $5\%$ as difficult. The average perplexity for tokens in each learning-difficulty level, obtained by Cross-Entropy Loss and the proposed MiLe Loss with $\gamma=1$, is shown in Figure 3. It can be seen that, compared with Cross-Entropy Loss, MiLe Loss results in an unnoticeable increase in perplexity for the easy tokens, while for the medium or the difficult tokens, MiLe Loss substantially reduces their perplexity with a noticeable decline. Given that easy tokens dominate the dataset, the overall increase in perplexity is expected. However, the substantial decline of perplexity for the medium or the difficult tokens further demonstrates the effectiveness of MiLe Loss in guiding language models to focus more on infrequent, difficult-to-learn but informative tokens and thereby mitigating the bias of learning difficulties during training.
$\gamma$ 0 0.5 1 2 5
PPL 5.473 5.467 5.492 5.608 6.317
Table 7: The perplexity (PPL) on the Pile validation set under different $\gamma$ values for MiLe Loss. Among them, $\gamma=0$ equals Cross-Entropy Loss.
Figure 3: The average perplexity (i.e., PPL) for tokens in different learning-difficulty levels.
6 Conclusions
In this paper, we present our observation of the bias of learning difficulties among tokens during language model pretraining, mainly caused by the inherent token imbalance in textual training data. We initially introduce Focal Loss as an attempt to mitigate the bias of learning difficulties. However, we find that considering the single probability of the ground-truth next token for assessing its learning difficulty is unreasonable, especially in cases with multiple valid next tokens. To tackle that, we propose MiLe Loss, which assesses the learning difficulty of a token by taking into account the global information entropy of the predicted probability distribution over the vocabulary. Extensive experiments demonstrate that, compared with both Cross-Entropy Loss and Focal Loss, the proposed MiLe Loss achieves superior performance for various downstream tasks in zero-shot and few-shot learning settings.
7 Limitations
In the proposed MiLe Loss, we scale the Cross-Entropy Loss based on information entropy to lead a generative language model to allocate more attention to difficult-to-learn tokens, which yields superior performance. Yet the effectiveness of MiLe Loss may be influenced by the quality of the training data. Specifically, as noisy data samples are generally outliers, the predicted probability distributions on them would typically exhibit high information entropy. Thus, too many noisy samples may make MiLe Loss amplify their corresponding loss weights too much, causing negative impacts on the model performance. We leave the investigation of how noisy data samples affect MiLe Loss to our future research.